The sum of two numbers is 40 and the difference is 16. What are the number?
Answer by macston(5194)

.
x and y are the two numbers
x-y=16
x=y+16 Use this to substitute for x
x+y=40 Substitute for x from above.
(y+16)+y=40 Subtract 16 from each side.
2y=24 Divide each side by 2.
y=12 ANSWER 1: One of the numbers is 12.
x=y+16=12+16=28 ANSWER 2: The other number is 28
CHECK:
x-y=16
28-12=16
16=16
x+y=40
28+12=40
40=40